# Thermal entanglement in conformal junctions

### v1

@@ 4.2 Logarithmic negativity
In this section, we compute the logarithmic negativity between $A$ and $B$ via replica trick, following Ref [51, 28]. To do so, we first introduce the Rényi negativity as
@@
with $\Delta_{b}$ the boundary dimension of $\mathcal{T}_{1}$ (or $\mathcal{T}_{2}$) with $n$ replicas. Eq. (4.23) follows from a general equivalence between the logarithmic negativity and the $1/2$-Rényi entropy [51] for bipartite systems (see also Ref. [52] for a related discussion on the ground state negativity).
@@
with $I_{1/2}(A,B)$ in Eq. (4.16). The relation above has occurred already in the literature in the context of quenches [53]. While (4.25) can be proven with elementary methods [^10] From the Schmidt decomposition of a pure state $\rho_{A\cup B}$ one can show that $\text{Tr}\left((\rho^{T_{B}}_{A\cup B})^{n_{e}}\right)=\text{Tr}\left((\rho_{A\cup B})^{n_{e}/2}\right)^{2}$, as explained in [51]. whenever $\rho_{A\cup B}$ is a pure state, it is not always valid, and simple counterexamples can be provided. In particular, it is not clearly a priori why (4.25) should hold exactly in our context, as the state we are considering is not pure. However, we believe that its validity strongly relies on the two-wire geometry considered above. For the sake of completeness, we mention that already for a three-wire geometry at zero temperature, analysed in detail in [13, 14], the mutual information between two wires does not obey (4.25). To conclude this section, we note that the above discussion can be generalized straightforwardly to $M>2$ wires. While the approach of Sec. 2 can be easily adapted, additional replica twist fields, acting non-trivially on subsets of wires (as in Eq. (4.5)), should be considered.
@@ 5 Lattice and numerics
We consider a lattice junction of free fermions, made by $M$ wires coupled at a single point via a conformal defect, that is a generalization of the model in Ref. [54, 55] where two wires were studied. We parametrize the point of the junctions via a semi-integer spatial index
@@
  • We perform a unitary transformation which diagonalizes $S$ while keeping the hopping term untouched and expressing the Hamiltonian in terms of decoupled nonphysical fields (see Ref. [13, 14] for details).
@@
at half-filling (see Ref. [56]), corresponding to the Fermi point $k_{F}=\pi/2$, and algebraic decay of correlations is present. While, as soon as $\beta$ is finite, the kernel $\mathcal{K}_{\beta}(x)$ decays exponentially to zero within a typical correlation length $\sim\beta$.
@@
In this limit, the thermal state of the junction is indistinguishable from the thermal state in the infinite volume limit, and the boundary does not play any role. While the latter property matches a simple physical intuition, it is not completely obvious, since non-trivial stationary states with long-range correlations between the wires can emerge in principle from the defect (see e.g. [57, 58, 59, 60, 61]).
@@ 5.1 Rényi entropies
To compute the Rényi entropy from the correlation functions, we follow Ref. [62, 63]. We first construct the correlation matrix $C$, reported in Eq. (5.6) for the half-line geometry, and we consider its restriction over a spatial subset $A$ which we denote as $C_{A}$. The Rényi entropy is thus expressed as
@@
parametrized by the scattering amplitude $\lambda\in[0,1]$. The system above has been studied in Refs. [54, 20, 64]. We first focus on an interval of length $\ell$ attached to the defect that belongs to the first wire: it corresponds to the sites $x\in[1/2,\dots,\ell-1/2]$ and $j=1$ in the Hamiltonian (5.3). To employ Eq. (4.8), we need to specify the values of scaling dimensions for the bulk/boundary twist field $\mathcal{T}^{1}$, defined in Eq. (4.5). In particular, we are interested in the limit $n\rightarrow 1$. Since the underlying CFT is the Dirac massless theory, whose central charge is $c=1$ [46], we have from Eq. (4.9)
@@
The boundary scaling dimension $\Delta_{b}$, is non-trivial and it depends explicitly on $\lambda$. Its value has been reported in [20], and it is related to the so-called effective central charge computed in [55]. In particular, the latter is defined as
@@
as physically expected, since the defect is not supposed to contribute to the entanglement in the geometry considered above. The prediction for $S_{n}$ has been reported in Eq. (4.15). We test it numerically and, for $L=30,\beta=30,\lambda=0.5$, we plot the results in Fig. 5. As expected, our formula Eq. (4.15), obtained in the limit $L\rightarrow\infty$, correctly describes the growth of the entropy until $\ell\simeq L$, where finite-size effects become relevant. In addition, we mention the presence of oscillations around the analytical predictions. While their characterization is beyond the scope of this work, we conjecture that they are related to the lattice corrections to the entropy discussed in detail in Ref. [56].
@@ 5.2 Negativity
In this section, we review the methods to compute negativity in the lattice, following closely Ref. [13]. We employ the notion of fermionic negativity, defined in [65, 66, 67, 31, 68, 69], and related to a notion of time-reversal transposition, that is naturally suited for free fermions.
@@
$\mathcal{E}_{n_{e}}\equiv\log\text{Tr}\left(|\rho_{A\cup B}|^{n_{e}}\right)=\text{Tr}\log\left(\left(\frac{1+\Gamma^{\times}_{A\cup B}}{2}\right)^{n_{e}/2}+\left(\frac{1-\Gamma^{\times}_{A\cup B}}{2}\right)^{n_{e}/2}\right)+$ (5.26)
@@
We consider the system of Sec, (5.1), namely we take two wires ($M=2$) and the scattering matrix $S$ in Eq. (5.18). We consider $A$ and $B$ as two intervals of length $\ell$ attached to the defect belonging to the first/second wire respectively. Specifically, the sites of $A$ are $x\in[1/2,\dots,\ell-1/2],j=1$ while the sites of $B$ are $x\in[1/2,\dots,\ell-1/2],j=2$. We need to specify the value of $\hat{\Delta}_{b}$ appearing in the analytical prediction Eq. (4.24). The latter gives the logarithmic prefactor of the ground state negativity that has been reported in [52]
@@ 6 Conclusions
Some questions are left open. For example, on the technical side, it would be interesting to characterize analytically multipoint-correlation functions of bulk and boundary twist fields. To the best of our knowledge, even in the case of free theories with many species and mixing boundary conditions (say, a non-trivial boundary scattering matrix) this task has not been accomplished. In principle, one could try to reformulate the boundary CFT problem as a chiral CFT via an unfolding map [46], where the latter depends on the boundary condition, and then adapt the techniques of Dixon et al. of orbifold CFT [70]. However, the main issue would be the presence of non-trivial monodromies induced by the twists, and a non-abelian Riemann-Hilbert problem would appear (see Ref. [71]). It is not clear to us if explicit analytical results can be obtained within this approach. Also, a deeper understanding of the boundary state ${}_{\beta}\bra{b}\sigma_{b}(0)$, introduced in Eq. (2.14), is preferable. In principle, since this is a state of the twisted Hamiltonian, it should be expressed as a linear combination of eigenstates of the latter. Still, it is not obvious whether it is non-normalizable (as it happens for the Ishibashi states [46]), and which eigenstates should appear in its expansion.
@@
Our methods have the potential to be adapted for addressing topological interfaces [72, 73, 74, 75]. In such instances, defects exhibit purely transmissive behavior, leading to intriguing boundary effects governed by the Affleck-Ludwig term. These effects are linked to the universal structure constant, which we previously overlooked, as they manifest as $O(1)$ terms that were not thoroughly discussed. Similarly, one should be capable of modifying the formalism to characterize excited states, previously investigated in [76] for BCFT, in order to replicate the findings of [77, 64]. It would be of interest to explore whether relevant defects can be examined similarly. In such cases, the boundary conditions lack scale invariance, introducing a non-trivial renormalization group flow. While perturbation theory suffices for short distances, a systematic approach to addressing large distances remains unclear (see Ref. [6]).
@@
Another possible application is the effect of symmetry-breaking boundary conditions in CFT for thermal states, extending the results of Ref. [47] of the ground state. We expect an exponential decay or saturation of the string-order parameter, as a consequence of a finite correlation length, and the results of Sec. 2 should apply directly.
@@
An interesting question, that remains so far unsolved, is the behavior of the entropy of a finite interval attached to the defect after a global quench. One expects that, for large enough time, the stationary state is reached, and, according to the results of this work, together with the general analysis of [57] for lattice free fermions (see also Refs. [58]), the information of the defect should be lost. On the other end at short time a linear growth should be observed, and, from the results of [20, 19], the rate should be proportional to $c_{\text{eff}}+c$. It is not clear to us how these two regimes can be interpolated. A scenario has been recently proposed in Ref. [78], where the authors conjecture that the entropy should be proportional to $c_{\text{eff}}+c$ at large times: to us, this conjecture looks unphysical, and incompatible with the results of this and previous works as well. We believe that, by adapting the results of [20], and performing a careful analysis of a two-point function of twist fields, one should be able to tackle the problem.
@@ Appendix A The boundary scaling dimension of twist fields
In this appendix, we give a detailed description of the boundary scaling dimensions of the twist fields, following standard techniques of BCFT [46, 79]. In particular, we first relate them to the ground state energy of a specific finite-size Hamiltonian. Then, via modularity, we show that it corresponds to a thermal free energy density, where non-trivial boundary conditions are kept along the (Euclidean) time direction. Let us consider the geometry
@@
via the transformation $z=\exp\left(-\frac{i\pi\zeta}{L}\right)$. The boundary conditions at $\text{Re}(\zeta)=\mp L/2$ are $b,b^{\prime}$ respectively. The resulting geometry gives the zero-temperature phase of a finite-size Hamiltonian $H_{bb^{\prime}}$ defined on a segment of length $L$. One can show via BCFT techniques that $H_{bb^{\prime}}$ is described by a chiral CFT [46, 43], and it holds
@@
Together with (A.5), Eq. (A.11) is the main result of this appendix, and it gives a physical interpretation of the boundary scaling dimension $\Delta_{b}$ of the twist field $\sigma$. We mention that detailed calculations of the overlap (A.8) have been performed in [9, 49, 12, 13, 14] for free theories to characterize the ground state entanglement of conformal junctions. However, their relation with the boundary twist fields is, up to our knowledge, new, albeit it follows directly from standard BCFT arguments.

### v2

@@ 4.2 Logarithmic negativity
In this section, we compute the logarithmic negativity between $A$ and $B$ via replica trick, following Ref [cct-12, cct-14]. To do so, we first introduce the Rényi negativity as
@@
with $\Delta_{b}$ the boundary dimension of $\mathcal{T}_{1}$ (or $\mathcal{T}_{2}$) with $n$ replicas. Eq. (4.23) follows from a general equivalence between the logarithmic negativity and the $1/2$-Rényi entropy [cct-12] for bipartite systems (see also Ref. [ge-20] for a related discussion on the ground state negativity).
@@
with $I_{1/2}(A,B)$ in Eq. (4.16). The relation above has occurred already in the literature in the context of quenches [ac-19]. While (4.25) can be proven with elementary methods [^10] From the Schmidt decomposition of a pure state $\rho_{A\cup B}$ one can show that $\text{Tr}\left((\rho^{T_{B}}_{A\cup B})^{n_{e}}\right)=\text{Tr}\left((\rho_{A\cup B})^{n_{e}/2}\right)^{2}$, as explained in [cct-12]. whenever $\rho_{A\cup B}$ is a pure state, it is not always valid, and simple counterexamples can be provided. In particular, it is not clearly a priori why (4.25) should hold exactly in our context, as the state we are considering is not pure. However, we believe that its validity strongly relies on the two-wire geometry considered above. For the sake of completeness, we mention that already for a three-wire geometry at zero temperature, analysed in detail in [cmc-22, cmc-22a], the mutual information between two wires does not obey (4.25). To conclude this section, we note that the above discussion can be generalized straightforwardly to $M>2$ wires. While the approach of Sec. 2 can be easily adapted, additional replica twist fields, acting non-trivially on subsets of wires (as in Eq. (4.5)), should be considered.
@@ 5 Lattice and numerics
We consider a lattice junction of free fermions, made by $M$ wires coupled at a single point via a conformal defect, that is a generalization of the model in Ref. [ep-12, pe-12] where two wires were studied. We parametrize the point of the junctions via a semi-integer spatial index
@@
  • We perform a unitary transformation which diagonalizes $S$ while keeping the hopping term untouched and expressing the Hamiltonian in terms of decoupled nonphysical fields (see Ref. [cmc-22, cmc-22a] for details).
@@
at half-filling (see Ref. [fc-11]), corresponding to the Fermi point $k_{F}=\pi/2$, and algebraic decay of correlations is present. While, as soon as $\beta$ is finite, the kernel $\mathcal{K}_{\beta}(x)$ decays exponentially to zero within a typical correlation length $\sim\beta$.
@@
In this limit, the thermal state of the junction is indistinguishable from the thermal state in the infinite volume limit, and the boundary does not play any role. While the latter property matches a simple physical intuition, it is not completely obvious, since non-trivial stationary states with long-range correlations between the wires can emerge in principle from the defect (see e.g. [bf-16, lsp-19, gls-23, fg-23, csrc-23]).
@@ 5.1 Rényi entropies
To compute the Rényi entropy from the correlation functions, we follow Ref. [Peschel-03, Peschel-09]. We first construct the correlation matrix $C$, reported in Eq. (5.6) for the half-line geometry, and we consider its restriction over a spatial subset $A$ which we denote as $C_{A}$. The Rényi entropy is thus expressed as
@@
parametrized by the scattering amplitude $\lambda\in[0,1]$. The system above has been studied in Refs. [ep-12, ce-23, ce-23a]. We first focus on an interval of length $\ell$ attached to the defect that belongs to the first wire: it corresponds to the sites $x\in[1/2,\dots,\ell-1/2]$ and $j=1$ in the Hamiltonian (5.3). To employ Eq. (4.8), we need to specify the values of scaling dimensions for the bulk/boundary twist field $\mathcal{T}^{1}$, defined in Eq. (4.5). In particular, we are interested in the limit $n\rightarrow 1$. Since the underlying CFT is the Dirac massless theory, whose central charge is $c=1$ [dms-97], we have from Eq. (4.9)
@@
The boundary scaling dimension $\Delta_{b}$, is non-trivial and it depends explicitly on $\lambda$. Its value has been reported in [ce-23], and it is related to the so-called effective central charge computed in [pe-12]. In particular, the latter is defined as
@@
as physically expected, since the defect is not supposed to contribute to the entanglement in the geometry considered above. The prediction for $S_{n}$ has been reported in Eq. (4.15). We test it numerically and, for $L=30,\beta=30,\lambda=0.5$, we plot the results in Fig. 5. As expected, our formula Eq. (4.15), obtained in the limit $L\rightarrow\infty$, correctly describes the growth of the entropy until $\ell\simeq L$, where finite-size effects become relevant. In addition, we mention the presence of oscillations around the analytical predictions. While their characterization is beyond the scope of this work, we conjecture that they are related to the lattice corrections to the entropy discussed in detail in Ref. [fc-11].
@@ 5.2 Negativity
In this section, we review the methods to compute negativity in the lattice, following closely Ref. [cmc-22]. We employ the notion of fermionic negativity, defined in [rac-16b, ssr-17, srrc-19, sr-19, sr-19a, srgr-18], and related to a notion of time-reversal transposition, that is naturally suited for free fermions.
@@
$\mathcal{E}_{n_{e}}\equiv\log\text{Tr}\left(|\rho_{A\cup B}|^{n_{e}}\right)=\text{Tr}\log\left(\left(\frac{1+\Gamma^{\times}_{A\cup B}}{2}\right)^{n_{e}/2}+\left(\frac{1-\Gamma^{\times}_{A\cup B}}{2}\right)^{n_{e}/2}\right)+$
@@
We consider the system of Sec, (5.1), namely we take two wires ($M=2$) and the scattering matrix $S$ in Eq. (5.18). We consider $A$ and $B$ as two intervals of length $\ell$ attached to the defect belonging to the first/second wire respectively. Specifically, the sites of $A$ are $x\in[1/2,\dots,\ell-1/2],j=1$ while the sites of $B$ are $x\in[1/2,\dots,\ell-1/2],j=2$. We need to specify the value of $\hat{\Delta}_{b}$ appearing in the analytical prediction Eq. (4.24). The latter gives the logarithmic prefactor of the ground state negativity that has been reported in [ge-20]
@@ 6 Conclusions
Some questions are left open. For example, on the technical side, it would be interesting to characterize analytically multipoint-correlation functions of bulk and boundary twist fields. To the best of our knowledge, even in the case of free theories with many species and mixing boundary conditions (say, a non-trivial boundary scattering matrix) this task has not been accomplished. In principle, one could try to reformulate the boundary CFT problem as a chiral CFT via an unfolding map [dms-97], where the latter depends on the boundary condition, and then adapt the techniques of Dixon et al. of orbifold CFT [dfms-87]. However, the main issue would be the presence of non-trivial monodromies induced by the twists, and a non-abelian Riemann-Hilbert problem would appear (see Ref. [mf-23]). It is not clear to us if explicit analytical results can be obtained within this approach. Also, a deeper understanding of the boundary state ${}_{\beta}\bra{b}\sigma_{b}(0)$, introduced in Eq. (2.14), is preferable. In principle, since this is a state of the twisted Hamiltonian, it should be expressed as a linear combination of eigenstates of the latter. Still, it is not obvious whether it is non-normalizable (as it happens for the Ishibashi states [dms-97]), and which eigenstates should appear in its expansion.
@@
Our methods have the potential to be adapted for addressing topological interfaces [gm-16, rs-22, rpr-22, hfsr-23]. In such instances, defects exhibit purely transmissive behavior, leading to intriguing boundary effects governed by the Affleck-Ludwig term. These effects are linked to the universal structure constant, which we previously overlooked, as they manifest as $O(1)$ terms that were not thoroughly discussed. Similarly, one should be capable of modifying the formalism to characterize excited states, previously investigated in [txas-13] for BCFT, in order to replicate the findings of [sths-22, ce-23a]. It would be of interest to explore whether relevant defects can be examined similarly. In such cases, the boundary conditions lack scale invariance, introducing a non-trivial renormalization group flow. While perturbation theory suffices for short distances, a systematic approach to addressing large distances remains unclear (see Ref. [kss-21]).
@@
Another possible application is the effect of symmetry-breaking boundary conditions in CFT for thermal states, extending the results of Ref. [bcp-23] of the ground state. We expect an exponential decay or saturation of the string-order parameter, as a consequence of a finite correlation length, and the results of Sec. 2 should apply directly.
@@
An interesting question, that remains so far unsolved, is the behavior of the entropy of a finite interval attached to the defect after a global quench. One expects that, for large enough time, the stationary state is reached, and, according to the results of this work, together with the general analysis of [bf-16] for lattice free fermions (see also Refs. [lsp-19]), the information of the defect should be lost. On the other end at short time a linear growth should be observed, and, from the results of [ce-23, wwr-17], the rate should be proportional to $c_{\text{eff}}+c$. It is not clear to us how these two regimes can be interpolated. A scenario has been recently proposed in Ref. [kkosw-23], where the authors conjecture that the entropy should be proportional to $c_{\text{eff}}+c$ at large times: to us, this conjecture looks unphysical, and incompatible with the results of this and previous works as well. We believe that, by adapting the results of [ce-23], and performing a careful analysis of a two-point function of twist fields, one should be able to tackle the problem.
@@ Appendix A The boundary scaling dimension of twist fields
In this appendix, we give a detailed description of the boundary scaling dimensions of the twist fields, following standard techniques of BCFT [dms-97, Recknagel:2013uja]. In particular, we first relate them to the ground state energy of a specific finite-size Hamiltonian. Then, via modularity, we show that it corresponds to a thermal free energy density, where non-trivial boundary conditions are kept along the (Euclidean) time direction. Let us consider the geometry
@@
via the transformation $z=\exp\left(-\frac{i\pi\zeta}{L}\right)$. The boundary conditions at $\text{Re}(\zeta)=\mp L/2$ are $b,b^{\prime}$ respectively. The resulting geometry gives the zero-temperature phase of a finite-size Hamiltonian $H_{bb^{\prime}}$ defined on a segment of length $L$. One can show via BCFT techniques that $H_{bb^{\prime}}$ is described by a chiral CFT [dms-97, Cardy-04], and it holds
@@
Together with (A.5), Eq. (A.11) is the main result of this appendix, and it gives a physical interpretation of the boundary scaling dimension $\Delta_{b}$ of the twist field $\sigma$. We mention that detailed calculations of the overlap (A.8) have been performed in [ss-08, bb-15, gm-17, cmc-22, cmc-22a] for free theories to characterize the ground state entanglement of conformal junctions. However, their relation with the boundary twist fields is, up to our knowledge, new, albeit it follows directly from standard BCFT arguments.